OpenStudy (anonymous):

Convert this equation from standard form to slope-intercept form... 3x-5y=25 .

hartnn (hartnn):

slope intercept form is y=mx+c so try to isolate y from 3x-5y=25 can u ?

OpenStudy (anonymous):

here slope = -a/b = -3/-5 = 3/5 y intercept = -c/b =25/-5 = -5 y = 3/5 * x - 5

OpenStudy (anonymous):

Thanks ! I got the same answer . (:

OpenStudy (anonymous):

Welcome.....